OpenStudy (anonymous):

What value of x makes this equation true? 2(x-5)-7=2 A. 19/2 B. -5 C. 7 D. -15/2

OpenStudy (nurali):

2(x-5)-7=2 Add both side by 7 2(x-5)-7+7=2+7 2(x-5)=9 Divide by 2 both side (x-5)=9/2 Add both side by 5 x-5+5=9/2+5 x=9/2+5 x=19/2